A 885 mL NaCl solution is diluted to a volume of 1.04 L and a concentration of 6.00 mol/L . What was the initial concentration?
1
Expert's answer
2019-04-30T04:31:30-0400

n=Cм х V Cм=n/V

n (NaCl)=6 x 1.04 = 6.24 mol

Cм=6.24/0.885=7.05 mol